a company produces mosquito repellents that have a period of effectiveness that is approximately normally distributed with a mean of 300 hours and a standard deviation of 30 hours. A sample of 40 repellents was taken and the average period of effectiveness is computed as 290 hours. Test the hypothesis that µ = 300 against the alternative that µ ≠ 300


1
Expert's answer
2022-06-08T18:05:21-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=300"

"H_1:\mu\not=300"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

The rejection region for this two-tailed test is "R = \{z:|z|> 1.96\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{290-300}{30/\sqrt{40}}=-2.1082"


6. Since it is observed that "|z|=2.1082>1.96=z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed is "p=2P(Z<-2.1082)=0.035014," and since "p=0.035014<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is different than 300, at the "\alpha = 0.05" significance level.